Suppose that a monopolist is producing a level of output such that AVC = $6, AFC = $4, P = $8, MR = $10, and MC = $6. Based on this information, the firm is realizing:
a loss which could be reduced by reducing price and increasing output
a profit which could be increased by reducing price and increasing output
a loss which could be reduced by increasing price and reducing output
a profit which could be increased by increasing price and reducing output
1
Expert's answer
2014-12-02T13:22:32-0500
If a monopolist is producing a level of output such that AVC = $6, AFC = $4, P = $8, MR = $10, and MC = $6, then its total profit is:
TP = (P - ATC)*Q, and as ATC = AVC + AFC = 6 + 4 = $10 > P = $8, so the firm faces losses, and as MR > MC (10 > 6), then a monopolist should decrease price increasing the output.
So, based on this information, the firm is realizing:
a) a loss which could be reduced by reducing price and increasing output